OpenStudy (anonymous):

What method would I use to integrate this? sqrt(4x^2+25x^4)

OpenStudy (anonymous):

I'm thinking trig sub but I can't figure out how to weasel that in.

OpenStudy (dumbcow):

\[x = \frac{4}{3} \tan u\] \[dx = \frac{4}{3} \sec^2 u du\]

OpenStudy (tkhunny):

Can you assume x > 0? \(\sqrt{x^{4} + 16x^{2}} = |x|\sqrt{x^{2} + 16}\) And a much simpler substitution may be possible.

OpenStudy (anonymous):

so your equation is\[\sqrt{9x^4+16x^2}\]you can pull out a common term of x^2 from both terms, and get\[ \pm x \sqrt{9x^2+16}\]integrate that and you're good

OpenStudy (tkhunny):

One cannot integrate "\(\pm\)". Use "+" for \(x\ge 0\) and "-" for \(x < 0\).

OpenStudy (anonymous):

I do believe that the x is positive, between 0 and 1. I see what you're saying. So I can take the x^2 out of the radical completely that way and then substitute tangeant inside it?

OpenStudy (tkhunny):

Tangent? I guess if you really want. I'd try u^2 = 9x^2 + 16

OpenStudy (dumbcow):

yeah do that ^^

OpenStudy (anonymous):

I'm trying to do that but I'm usually awful at visualizing u substitution for some reason. :P So u^2 = 9x^2+16. Is du=18x? So you get (sqrt(u) du)/18?

OpenStudy (dumbcow):

you need to differentiate the u^2 as well --> 2u du = 18x dx --> dx = u/9x dx

OpenStudy (anonymous):

Oh wow and that's it? All that's left to do is plug stuff back in to the u and simplify?

OpenStudy (dumbcow):

yeah after integrating

OpenStudy (anonymous):

Yeah I just realized that as I was writing. So to integrate that do you need to put 9x into terms of u?

OpenStudy (dumbcow):

normally yes, but in this case it will cancel out with the "x" you factored out of sqrt

OpenStudy (anonymous):

I'm pretty confused here. Where does the extra t come in to cancel?

OpenStudy (tkhunny):

"t"?

OpenStudy (anonymous):

I'm sorry I meant x

OpenStudy (tkhunny):

Simplify your life a little. \(u = 9x^{2} + 16\) \(du = 18x\;dx\) \(\int x\sqrt{9x^{2} + 16}\;dx = \dfrac{1}{18}\int 18x\sqrt{9x^{2} + 16}\;dx = \dfrac{1}{18}\int \sqrt{u}\;du = \dfrac{1}{18}\dfrac{u^{3/2}}{3/2} + C\)

OpenStudy (anonymous):

As I said before these substitutions kill me. XP That makes a bit more sense to me. I'll try to roll through it that way and see what happens. Turns out I thought I got the correct answer before but I didn't.

OpenStudy (anonymous):

so in the end you get: \[\frac{ 1 }{ 27 }(9t^2+16)^{3/2}\] right?

OpenStudy (anonymous):

Yes that seems to have done it for sure this time. Thanks a lot! I wish I could give medals to all of you!

OpenStudy (tkhunny):

Except for your missing Constant. You can always find the 1st derivative and check!